find three consecutive integers whose sum is -126

Answer by checkley71(8403)

LET X BE THE FIRST INTEGER.
X+1 IS THE SECOND INTEGER.
X+2 IS THE THIRD INTEGER.
X+X+1+X+2=-126
3X+3=-126
3X=-126-3
3X=-129
X=-129/3
X=-43 ANSWER.
-43+1=-42 ANSWER.
-43-+2=-41 ANSWER.
PROOF
-43-42-41=-126
-126=-126